OpenStudy (anonymous):

let me post my question as a comment plz. ^_^ Btw. i want to know how to solve this problem, i dont just want a straight up answer, i want to be guided to my answer. (P.s.) Thanks a ton! ^_^

OpenStudy (anonymous):

divide \[\sqrt{9x^2}/\sqrt{18y^2}\] I will give a medal. ^_^

OpenStudy (anonymous):

@AravindG @DLS @Callisto

OpenStudy (aravindg):

First use the fact that 

OpenStudy (anonymous):

yah, i have that much down. In fact, i already tried to get the answer, and i came out with \[x/\sqrt{2y}\]

OpenStudy (anonymous):

but that isn't an option in my choices.

OpenStudy (anonymous):

would it be \[s \sqrt{2}/2y\] ?

hartnn (hartnn):

do you mean, \(\dfrac{x}{\sqrt 2 y }\) ??

OpenStudy (anonymous):

yes, thats it

OpenStudy (anonymous):

yah, if there is a / than it means divide. i didnt know how to get a line like that

hartnn (hartnn):

what are your options?

hartnn (hartnn):

is \(\dfrac{\sqrt 2 x}{ 2 y }\) one of them ? because both are equivalent

OpenStudy (anonymous):

A \[3x/2y\] B \[\frac{ x \sqrt{2} }{ 2y }\] C \[\frac{ x \sqrt{3} }{ y }\] D 1/2xy

OpenStudy (anonymous):

those are my options

OpenStudy (anonymous):

i thought it was b, but i just wanted to confirm. Thanks @Loser66

OpenStudy (anonymous):

... his comment dissapeared.